While standing on the roof of a building, a child tosses a tennis ball with an initial speed of 12 m/s at an angle of 25° below the horizontal. The ball lands on the ground 3.3 s later. How tall, in meters, is the building?
Answer #1

Using 2nd kinematic equation in vertical direction:

h = Viy*t + (1/2)*ay*t^2

Viy = Initial velocity in vertical direction = V0*sin theta

Viy = 12*sin (-25 deg) = -5.07 m/sec (-ve sign because given that angle is below the horizontal)

ay = -g = -9.81 m/sec^2

t = time taken = 3.3 sec

So, Using these Values:

h = -5.07*3.3 - 0.5*9.81*3.3^2

h = -70.146 m = -70.1 m

Height of building = 70.1 m

Please Upvote.

Comment below if you have any query.